# Radial and Non-Radial Solution Structures for Quasilinear Hamilton–Jacobi–Bellman Equations in Bounded Settings

### v2

@@ 3.1.4 Legendre transform yielding the quasilinear elliptic PDE
The function $f(v):=\xi\cdot v+|v|^{\alpha}$ is strictly convex and coercive. The minimizer $v^{*}$ satisfies the first-order condition
$\xi+\alpha|v^{*}|^{\alpha-2}v^{*}=0,$
which gives $v^{*}=-\frac{1}{\alpha^{1/(\alpha-1)}}|\xi|^{1/(\alpha-1)}\frac{\xi}{|\xi|}$. Substituting back:
$f(v^{*})$ $=\xi\cdot\Big(-\frac{1}{\alpha^{1/(\alpha-1)}}|\xi|^{1/(\alpha-1)}\frac{\xi}{|\xi|}\Big)+\Big|\frac{1}{\alpha^{1/(\alpha-1)}}|\xi|^{1/(\alpha-1)}\frac{\xi}{|\xi|}\Big|^{\alpha}$
@@
4 Radial Symmetry of Solutions
When the domain and data possess radial symmetry, the solution inherits this symmetry. We establish this via the uniqueness result and derive the corresponding ordinary differential equation.
Theorem 4.1 (Radial symmetry).
Suppose $\Omega=B_{R}(0):=\{x\in\mathbb{R}^{N}:|x|<R\}$ is a ball, $h(x)=\tilde{h}(|x|)$ is radial, and $g$ is constant on $\partial\Omega$. Then the unique solution $V$ of (1.1) is radial: $V(x)=u(|x|)$ for some function $u:[0,R]\to\mathbb{R}$.
Proof.
For any orthogonal matrix $Q\in O(N)$ (i.e., $Q^{T}Q=I_{N}$ and $\det(Q)=\pm 1$), define $V_{Q}(x):=V(Qx)$. We verify that $V_{Q}$ also solves (1.1).
Step 1: PDE satisfied by $V_{Q}$. Compute the derivatives using the chain rule:
$\nabla V_{Q}(x)$ $=Q^{T}\nabla V(Qx),$
$\Delta V_{Q}(x)$ $=\text{tr}(D^{2}V_{Q}(x))=\text{tr}(Q^{T}D^{2}V(Qx)Q)=\text{tr}(D^{2}V(Qx))=\Delta V(Qx),$
where we used that $\text{tr}(Q^{T}AQ)=\text{tr}(A)$ for any matrix $A$. Therefore,
$-\frac{\sigma^{2}}{2}\Delta V_{Q}(x)+C_{\alpha}|\nabla V_{Q}(x)|^{p}-h(x)$ $=-\frac{\sigma^{2}}{2}\Delta V(Qx)+C_{\alpha}|Q^{T}\nabla V(Qx)|^{p}-h(Qx)$
$=-\frac{\sigma^{2}}{2}\Delta V(Qx)+C_{\alpha}|\nabla V(Qx)|^{p}-h(Qx)$
$=0,$
where we used $|Q^{T}\xi|=|\xi|$ for any $\xi\in\mathbb{R}^{N}$ (since $Q$ is orthogonal) and the fact that $h(Qx)=\tilde{h}(|Qx|)=\tilde{h}(|x|)=h(x)$ by radial symmetry.
Step 2: Boundary condition. On $\partial\Omega=\{x:|x|=R\}$, we have $|Qx|=|x|=R$, so $Qx\in\partial\Omega$. Since $g$ is constant on $\partial\Omega$, $V_{Q}(x)=V(Qx)=g$ for all $x\in\partial\Omega$.
Step 3: Uniqueness implies radial symmetry. By Proposition 2.10, the solution to (1.1) is unique. Since both $V$ and $V_{Q}$ solve the same problem, we have $V_{Q}\equiv V$, i.e., $V(Qx)=V(x)$ for all $Q\in O(N)$ and $x\in\overline{\Omega}$. This means $V$ is invariant under the orthogonal group $O(N)$, hence $V(x)=u(|x|)$ for some function $u:[0,R]\to\mathbb{R}$. ∎
Theorem 4.2 (Radial ODE).
The radial solution $u(r)$, where $r=|x|$, satisfies the ordinary differential equation
$\frac{\sigma^{2}}{2}\Big(u^{\prime\prime}(r)+\frac{N-1}{r}u^{\prime}(r)\Big)+C_{\alpha}\,\left|u^{\prime}(r)\right|^{p}+\tilde{h}(r)=0,\quad 0<r<R,$ (4.1)
with boundary conditions $u^{\prime}(0)=0$ and $u(R)=g$.
Proof.
For a radial function $V(x)=u(|x|)=u(r)$ with $r=|x|$, we compute the derivatives. Let $e_{r}:=x/|x|$ be the radial unit vector. Then
$\nabla V(x)=u^{\prime}(r)e_{r}=u^{\prime}(r)\frac{x}{r}.$
The Laplacian in radial coordinates is
$\Delta V(x)=u^{\prime\prime}(r)+\frac{N-1}{r}u^{\prime}(r).$
Moreover, $|\nabla V(x)|=|u^{\prime}(r)|$. Substituting into (1.1) yields
$\frac{\sigma^{2}}{2}\Big(u^{\prime\prime}(r)+\frac{N-1}{r}u^{\prime}(r)\Big)+C_{\alpha}\,|u^{\prime}(r)|^{p}+\tilde{h}(r)=0.$
The boundary conditions follow from: (i) $u^{\prime}(0)=0$ by regularity at the origin (otherwise $\nabla V$ would be singular), and (ii) $u(R)=g$ from the Dirichlet condition on $\partial B_{R}(0)$. ∎
Remark 4.3.
For the case $\alpha=2$ (corresponding to $p=2$), equation (4.1) becomes
$\frac{\sigma^{2}}{2}\Big(u^{\prime\prime}(r)+\frac{N-1}{r}u^{\prime}(r)\Big)+\frac{1}{4}(u^{\prime}(r))^{2}+\tilde{h}(r)=0,$
which can be solved explicitly in some cases using phase-plane methods.
5 Computational Implementation and Applications
The constructive nature of Theorem 1.1 makes it particularly suitable for numerical implementation. The monotone iteration scheme (2.11) provides an algorithmic approach that is both theoretically sound and computationally efficient.
5.1 Numerical algorithm
The following algorithm implements the monotone iteration:
  1. 1.
    Initialization: Solve $\Delta\phi=1$ in $\Omega$ with $\phi=0$ on $\partial\Omega$ to obtain the torsion function. Set $V^{(0)}=g+B\phi$ with $B\geq 2H/\sigma^{2}$.
  2. 2.
    Iteration: For $k=0,1,2,\ldots$, solve the linear PDE
@@ 5.1 Numerical algorithm
  3. 3.
    Convergence test: Stop when $\|V^{(k+1)}-V^{(k)}\|_{L^{\infty}(\Omega)}<\varepsilon$.
5.2 Software implementation
For the special case of radially symmetric problems ($\Omega=B_{R}(0)$ and $h$ radial), numerical codes implementing the above algorithm are available in our companion repository [2]. See
https://github.com/coveidragos/code_contrast_image/blob/main/contrast_image.py
for the denoised-restored contrast scheme, or
https://github.com/coveidragos/code_contrast_image/blob/main/contrast_image_rest_denoised.py
for the restored-denoised contrast scheme.
The codes solve the radial ODE (4.1) using finite difference methods with adaptive mesh refinement.
For general non-radial domains, the implementation requires finite element methods or spectral methods. Further details and guidance for extending the implementation to these settings can be found in [2, 3].
5.3 Applications
@@
6 Concluding Remarks
@@
Acknowledgments
The author thanks the anonymous referees for their careful reading and helpful suggestions that improved the presentation of this paper. The author is grateful to Professors P.L. Lions and J.M. Lasry for their pioneering work on Hamilton–Jacobi–Bellman equations, which provided the foundation for this study.
@@ Acknowledgments
No new data were created or analyzed in this study. The numerical implementation codes referenced in Section 6 are available at the public repository cited in [2].

### v3

@@ 4.1.3. Legendre transform yielding the quasilinear elliptic PDE
The function $f(v):=\xi\cdot v+|v|^{\alpha}$ is strictly convex and coercive. The minimizer $v^{\ast}$ satisfies the first-order condition
$\xi+\alpha|v^{\ast}|^{\alpha-2}v^{\ast}=0,$
which gives
$v^{\ast}=-\frac{1}{\alpha^{1/(\alpha-1)}}|\xi|^{1/(\alpha-1)}\frac{\xi}{|\xi|}.$
Substituting back:
$f(v^{\ast})$ $=\xi\cdot\Big(-\frac{1}{\alpha^{1/(\alpha-1)}}|\xi|^{1/(\alpha-1)}\frac{\xi}{|\xi|}\Big)+\Big|\frac{1}{\alpha^{1/(\alpha-1)}}|\xi|^{1/(\alpha-1)}\frac{\xi}{|\xi|}\Big|^{\alpha}$
@@
5. Numerical Analysis
This section details the numerical implementation of the quasilinear HJB framework and its applications in operational research and imaging. The constructive nature of Theorem 1.1 makes it particularly suitable for numerical implementation. The monotone iteration scheme provides an algorithmic approach that is both theoretically sound and computationally efficient.
5.1. Weighted linear monotone iteration and stochastic simulation
For the numerical approximation of the solution to the Hamilton–Jacobi–Bellman equation, we employ the weighted monotone iteration introduced by Lasry and Lions. This method is fully consistent with the analytical existence proof and is particularly robust on bounded domains. In addition, once the value function is computed, we construct the optimal feedback control and simulate the associated controlled diffusion.
The scheme constructs a decreasing sequence of functions $\{V^{(k)}\}_{k\geq 0}$ trapped between the ordered sub- and super-solutions
$V_{-}(x)=g,\qquad V_{+}(x)=g+B\phi(x),$
where $\phi$ is the torsion function solving $\Delta\phi=1$ in $\Omega$, $\phi=0$ on $\partial\Omega$, and $B\geq 2H/\sigma^{2}$ with $H=\max_{\overline{\Omega}}h$.
  1. (1)
    Preprocessing (torsion function and barriers). Compute the torsion function $\phi$ solving
    $\Delta\phi=1\quad\text{in }\Omega,\qquad\phi=0\quad\text{on }\partial\Omega.$
    Define the ordered barriers
    $V_{-}(x):=g,\qquad V_{+}(x):=g+B\phi(x),\quad B\geq\frac{2H}{\sigma^{2}},\quad H:=\max_{\overline{\Omega}}h.$
  2. (2)
    Initialization. Set the initial iterate
    $V^{(0)}:=V_{+}=g+B\phi.$
  3. (3)
    Weighted linear monotone iteration. For $k=0,1,2,\dots$ compute $V^{(k+1)}$ as the solution of
@@ 5.1. Weighted linear monotone iteration and stochastic simulation
    with Dirichlet boundary condition
    $V^{(k+1)}=g\quad\text{on }\partial\Omega.$
    The constant weight $\Lambda_{0}>0$ is chosen so that the nonlinear Hamiltonian is dominated by the linear term:
    $\Lambda_{0}\;\geq\;C_{\alpha}\,p\,M^{p-1}\,C_{\Omega},\qquad M:=\max\bigl\{\|\nabla V_{-}\|_{L^{\infty}},\|\nabla V_{+}\|_{L^{\infty}}\bigr\},$
    where $C_{\Omega}$ is a Poincaré-type constant for the domain.
  4. (4)
    Monotonicity and projection. After each linear solve, enforce the bounds
    $V_{-}\;\leq\;V^{(k+1)}\;\leq\;V_{+}\quad\text{in }\Omega,$
    by projecting pointwise onto the interval $V_{-},V_{+}]$. This guarantees that the sequence remains ordered and uniformly bounded.
  5. (5)
    Convergence criterion. Terminate the iteration when
    $\|V^{(k+1)}-V^{(k)}\|_{L^{\infty}(\Omega)}<\varepsilon,$
    for a prescribed tolerance $\varepsilon>0$. The resulting limit $V:=\lim_{k\rightarrow\infty}V^{(k)}$ is the numerical approximation of the value function.
  6. (6)
    Optimal feedback control. Once $V$ is computed, approximate its gradient $\nabla V$ on the grid and define the optimal feedback control $v^{\ast}=(v_{1}^{\ast},v_{2}^{\ast})$ by
    $v^{\ast}(x)=-\frac{1}{\alpha^{\,p-1}}\bigl|\nabla V(x)\bigr|^{p-2}\nabla V(x),\qquad p=\frac{\alpha}{\alpha-1},$
    with the convention $v^{\ast}(x)=0$ whenever $\nabla V(x)=0$ or $x\notin\Omega$.
  7. (7)
    Stochastic dynamics simulation. To illustrate the controlled dynamics, we simulate sample paths of the controlled diffusion
    $dX_{t}=v^{\ast}(X_{t})\,dt+\sigma\,dW_{t},$
    starting from an initial state $X_{0}=x_{0}\in\Omega$. On a time grid $t_{n}=n\Delta t$, $n=0,\dots,N$, we use the Euler–Maruyama scheme
    $X_{n+1}=X_{n}+v^{\ast}(X_{n})\,\Delta t+\sigma\sqrt{\Delta t}\,\xi_{n},$
    where $(\xi_{n})_{n\geq 0}$ are independent standard Gaussian vectors in $\mathbb{R}^{2}$. The simulation is stopped when $X_{n}$ approaches the boundary of $\Omega$ (e.g., when the distance to $\partial\Omega$ falls below a prescribed margin), and the resulting trajectory is used to visualize the inventory dynamics under the optimal policy.
The operator on the left-hand side of (5.1) is coercive and order-preserving, and the right-hand side is monotone in $V^{(k)}$. Therefore the sequence $\{V^{(k)}\}$ is decreasing and converges to the unique classical solution of the HJB equation. This scheme is stable, monotone, and preserves the concavity of the value function in the quadratic case $\alpha=2$ discovered in [2], while the stochastic simulation step provides a natural Monte Carlo illustration of the optimal controlled dynamics.
5.2. Software implementation (radially symmetric case)
For the special case of radially symmetric problems ($\Omega=B_{R}(0)$ and $h$ radial), numerical codes implementing the above algorithm are available in our companion repository [4].
5.3. Applications to Production Planning and Image Enhancement
@@
We present a comprehensive numerical validation of the monotone iteration scheme. The implementation, developed in Python using high-precision sparse solvers, demonstrates the versatility of the quasilinear HJB framework in both operational research and computer vision. The codes solve the 2D quasilinear HJB PDE (1.1) directly on high-resolution grids (e.g., $241\times 241$) using finite difference methods with efficient sparse solvers. This robust framework handles both the radial benchmark cases (1.2) and general non-radial domains, as extensively documented in [4, 3].
5.3.1. Stochastic Production Planning Dynamics
We evaluate the optimal production policy for a system governed by $\alpha=1.5$ (leading to a cubic gradient penalty $p=3$) on an elliptical inventory domain $\Omega$. With parameters $\sigma=0.3$ and boundary cost $g=0.5$, the Policy Iteration solver converges rapidly to a tolerance of $10^{-6}$ in approximately 9 iterations. As illustrated in Figure 1, the value function $V(y)$ is strictly positive and concave, characterizing the minimal expected cost associated with varying inventory levels. The values range from $0.5$ at the boundary to a peak of approximately $1.67$ at the center, establishing a stable feedback production rate that drives the inventory toward the optimal set-point.
Figure 1. Numerical results for the stochastic production planning problem with $\alpha=1.5$ and $h(y)=|y|^{p}$. Left: Value function $V(y)$. Center/Right: Optimal control components and simulated inventory trajectory.
Figure 2. Numerical approximation of the value function $V(y)$ on the elliptical domain $\Omega=\{y:(y_{1}/1.5)^{2}+y_{2}^{2}<1\}$ for $\alpha=1.5$, $h(y)=|y|^{p}$, $\sigma=0.3$, and boundary data $g=0.5$. The surface rises monotonically from $V=0.5$ on $\partial\Omega$ to a maximum of approximately $1.67$ at the center, confirming the positivity result of Theorem 1.1. The Policy Iteration scheme converges in 9 steps to a tolerance of $10^{-6}$.
Our Python code implementing the monotone iteration scheme used in the proofs of our results for the general quasilinear HJB elliptic equation on bounded domains, and which was used to generate the plots in this paper, can be found in [8].
5.3.2. Contrast Enhancement via Nonlinear Hamilton–Jacobi–Bellman Dynamics
Contrast enhancement is a fundamental operation in image processing, with direct impact on visual perception, feature extraction, and downstream computer vision tasks. Classical approaches such as histogram equalization or total variation (TV) filtering operate either through global redistribution of intensities or local smoothing mechanisms. While effective in specific scenarios, these methods lack a principled mechanism for controlling the enhancement strength in relation to the underlying image geometry.
We propose a nonlinear contrast enhancement framework derived from the stationary Hamilton–Jacobi–Bellman (HJB) equation
$\frac{\sigma^{2}}{2}\Delta V(y)+C_{\alpha}\,|\nabla V(y)|^{p}-h(y)=0,\qquad p=\frac{\alpha}{\alpha-1},$
where $V$ is the value function defined on the chromaticity plane, $h(y)$ is a user-defined potential, and $\alpha>1$ is the key parameter controlling the nonlinearity of the system. The enhanced image is obtained by advecting pixel values along the optimal drift field
$b(y)=A_{\alpha}\,|\nabla V(y)|^{\frac{1}{\alpha-1}}\frac{\nabla V(y)}{|\nabla V(y)|},\qquad A_{\alpha}=\left(\frac{1}{\alpha}\right)^{\frac{1}{\alpha-1}}.$
Role of the Nonlinearity Parameter $\alpha$
A central finding of our study is that the parameter $\alpha$ is the dominant factor governing the strength of contrast enhancement. When $\alpha$ approaches $1^{+}$, the exponent $p=\alpha/(\alpha-1)$ becomes large, amplifying the nonlinear gradient term and producing strong geometric enhancement. Conversely, larger values of $\alpha$ reduce the nonlinearity, yielding milder transformations.
This behavior is consistent with the theoretical structure of HJB equations: the term $|\nabla V|^{p}$ acts as a geometric amplifier of local intensity variations, and $\alpha$ directly modulates the sensitivity of the system to these variations. In contrast, the potential $h(y)$ influences only the forcing term and has a secondary effect on the global contrast dynamics.
Experimental Evaluation
We evaluate the proposed method on a representative RGB image and compare it against classical histogram equalization and a post-processed HJB+TV variant. Contrast quality is assessed using three widely adopted metrics: the Enhancement Measure Estimation (EME), the Tenengrad sharpness index, and the standard deviation of luminance.
Table 1 summarizes the results for a sweep of $\alpha$ values in the interval $[1.02,2.0]$.
Table 1. Contrast metrics for various values of $\alpha$. The proposed HJB method achieves its strongest enhancement for $\alpha\in[1.02,1.10]$, significantly outperforming classical methods.
Method EME Tenengrad StdDev
Original 5.59 106.63 70.91
Histogram Equalization 4.07 49.12 84.39
HJB ($\alpha=1.02$$1.10$) 7.97 187.16 85.53
HJB+TV ($\alpha=1.02$$1.10$) 10.06 148.72 85.28
HJB ($\alpha=1.15$) 9.28 153.42 79.87
HJB+TV ($\alpha=1.15$) 7.05 117.52 79.62
HJB ($\alpha=1.20$) 6.84 123.08 74.19
HJB+TV ($\alpha=1.20$) 4.86 90.57 73.97
HJB ($\alpha=1.30$) 6.01 112.05 72.13
HJB+TV ($\alpha=1.30$) 4.30 81.86 71.93
HJB ($\alpha=1.50$) 5.81 109.15 71.53
HJB+TV ($\alpha=1.50$) 4.18 79.62 71.33
HJB ($\alpha=2.0$) 5.72 107.67 71.30
HJB+TV ($\alpha=2.0$) 4.12 78.47 71.10
The results reveal a clear and consistent pattern:
  • For $\alpha\in[1.02,1.10]$, the HJB method produces the strongest enhancement, achieving EME = 7.97 and Tenengrad = 187.16, while the HJB+TV variant reaches EME = 10.06. This represents a substantial improvement over histogram equalization.
  • The enhancement strength decreases gradually as $\alpha$ increases. The parameter $\alpha$ acts as a global contrast-control knob: small values produce aggressive sharpening, while larger values yield smoother, more conservative enhancement.
  • For $\alpha\geq 1.3$, the method behaves similarly to a mild diffusion process, producing results close to the original image but still improving edge sharpness moderately.
Figure 3. Comparison between the original image, histogram equalization, the proposed HJB 2D PDE method ($\alpha=1.15$), with $h(y)=|y|^{p}$, and the HJB 2D + TV variant. The composite figure illustrates the strong contrast enhancement achieved by the HJB approach, which significantly outperforms classical methods both visually and quantitatively.
Discussion
These findings demonstrate that the proposed HJB-based framework provides a mathematically rigorous and highly tunable mechanism for contrast enhancement. Unlike classical methods that lack a principled control parameter, the HJB formulation offers a single scalar $\alpha$ that continuously interpolates between strong geometric enhancement and gentle diffusive smoothing. This versatility, combined with superior performance in high-contrast regimes, establishes the HJB approach as a significant contribution to the field of variational image processing.
6. Conclusions and Future Directions
@@
No new data were created or analyzed in this study.
@@
Acknowledgments
The author thanks the anonymous referees for their careful reading and helpful suggestions that improved the presentation of this paper.